OpenStudy (anonymous):

Find a polynomial for the perimeter and for the area. There is a box with d+2 on top and on the right side there is a d. Need the perimeter, and the area.

OpenStudy (anonymous):

so lets say a,b,c are the sides than a=b=d+2 and c=d+2?

OpenStudy (anonymous):

i wrote d+2 twice :)

OpenStudy (anonymous):

the box has area a*b*c

OpenStudy (anonymous):

and perimeter: 2ab+2ac+2bc

OpenStudy (anonymous):

I honestly don't know how to do it this is what the question was.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

hey can I ask you something?

OpenStudy (anonymous):

you have to put d+2 instead of a,b and d instead of c

OpenStudy (anonymous):

so area will be (d+2)^2*d=d^3+4d^2+4d

OpenStudy (anonymous):

perimeter is 2(d+2)(d+2) +4d(d+2)

OpenStudy (anonymous):

if I understood the question correctly, the phrasing is pretty bad

OpenStudy (anonymous):

Is that after simplifing the answer?

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

both the perimeter and the area have to be simpfly not factor.

OpenStudy (anonymous):

2(d+2)(d+2) +4d(d+2)=2d^2+8d+8+4d^2+8d=6d^2+16d+8

OpenStudy (anonymous):

thank you can you do my other sample i have too?

OpenStudy (anonymous):

can you do it? I guess this is some kind of homework :-) I can help but I wont solve it for you.

OpenStudy (anonymous):

\[\sqrt[20]{(-9)^{20}}\]

OpenStudy (anonymous):

ok. sounds good to me.

OpenStudy (anonymous):

is that 30sqrt(-9)^20 ? not really visible

OpenStudy (anonymous):

20sqrt(-9)^20

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

what is ur guess?

OpenStudy (anonymous):

6.088

OpenStudy (anonymous):

how did you get that? :)

OpenStudy (anonymous):

i times -9 20 times and then i divied it by 20

OpenStudy (anonymous):

oh I see, but here this is not multiplication or division

OpenStudy (anonymous):

what does 2^4 mean?

OpenStudy (anonymous):

the 4 is the root of 2?

OpenStudy (anonymous):

would be 8

OpenStudy (anonymous):

nope, 2 to the power of 4 means that you take 2*2*2*2 (4times)=16

OpenStudy (anonymous):

so (-9)^20=-9*-9*-9..... (20 times)

OpenStudy (anonymous):

now what does \[\sqrt{?}\] mean?

OpenStudy (anonymous):

it is the inverse operation so lets say \[\sqrt[3]{8}\]

OpenStudy (anonymous):

that means the number that you multiply by itself 3 times and you will get 8

OpenStudy (anonymous):

when you multiply -9 20 times you don't get a good answer

OpenStudy (anonymous):

you get a horrid number, really big one

OpenStudy (anonymous):

but now we have 20sqrt(-9)^20

OpenStudy (anonymous):

what is the number that you multiply by itself 20 times to get (-9)^20?

OpenStudy (anonymous):

(-9)^20 means -9 multiplied by itself 20 times

OpenStudy (anonymous):

i know what i get when u multiply -9. 20 times. u get 282042115992

OpenStudy (anonymous):

I guess you counted that in your head :-)

OpenStudy (anonymous):

no used my phone.

OpenStudy (anonymous):

you do not need the exact value to solve the problem

OpenStudy (anonymous):

the answer is: what is the number that you multiply by itself 20 times to get (-9)^20? (-9)^20 means -9 multiplied by itself 20 times it is -9

OpenStudy (anonymous):

but that is not the good answer as the correct one is 9 can you see why?

OpenStudy (anonymous):

by the way, I suggest not using calculator for studying maths, unless there is some horrid calculation.